OpenStudy (anonymous):

(2x+5y) Exponent of 2 on the outside is talking about binomials i know how to combine but i dont know what it is asking here

OpenStudy (anonymous):

What is the question asking you to do with the binomial? Simplify?

OpenStudy (anonymous):

its telling me to consider the special product

OpenStudy (anonymous):

but it asking for a simplified version of that

OpenStudy (dumbcow):

(a+b)^2 = a^2 +2ab + b^2

OpenStudy (anonymous):

i dont get that?

OpenStudy (dumbcow):

that is the general form for a squared binomial (2x+5y)^2 = (2x)^2 + 2(2x*5y) + (5y)^2

OpenStudy (anonymous):

so with that equation I got4x^2+40xy+25y^2

OpenStudy (anonymous):

It was wrong

OpenStudy (dumbcow):

2*2*5 = 20

OpenStudy (anonymous):

Yay I get it thanks!